# Beyond Face Swapping: A Diffusion-Based Digital Human Benchmark for Multimodal Deepfake Detection

### v1

@@ 3.3. Dataset Organization
Forged videos and audio can form four combinations: (1) real video–real audio (RV-RA), (2) real video–fake audio (RV-FA), (3) fake video–real audio (FV-RA), and (4) fake video–fake audio (FV-FA). We exclude RV-FA for two reasons: First, our focus is on video-based facial forgery detection, whereas RV-FA does not involve human facial manipulation. For example, DFDC (dolhansky2020deepfakedetectionchallengedfdc, ) labels RV-FA samples as fake, which is inherently unfair to vision-only detectors. Second, Audio-only manipulations without visual changes are often obvious and pose limited detection challenges. For example, FakeAVCeleb (khalid2022fakeavcelebnovelaudiovideomultimodal, ) contains audio-video synchronization delays.
We then introduce three combinations and the corresponding methods adopted in this study, with detailed explanations in subsequent use cases:
@@
Fake Video–Real Audio (FV-RA) This category contains fake videos generated by synthesizing visual content conditioned on real audio. We first extract and convert audio from real videos into WAV format. RetinaFace is then used to detect and crop representative facial frames. Using five digital human generation methods: Sonic (ji2024sonicshiftingfocusglobal, ), Hallo1 (xu2024hallohierarchicalaudiodrivenvisual, ), Hallo2 (cui2024hallo2longdurationhighresolutionaudiodriven, ), Echomimic (chen2024echomimiclifelikeaudiodrivenportrait, ), and V-Express (wang2024vexpressconditionaldropoutprogressive, ), we generate a total of 25,000 forged videos. Modern attackers can acquire the speech and facial data of a target to conduct highly realistic impersonation attacks, surpassing the threats of traditional face-swapping methods.
Fake Video–Fake Audio (FV-FA). This category encompasses both forged audio and video. We employed a state-of-the-art voice cloning method, CosyVoice 2 (du2024cosyvoice2scalablestreaming, ). Specifically, we first use a large language model to generate processed text. By providing real audio-text pairs, the CosyVoice 2 model synthesizes fake audio mimicking the target speaker’s voice characteristics. We then generate 25,000 forged videos based on this fake audio using algorithms like Sonic, Hallo1, Hallo2, and Echomimic. This subset helps researchers study advanced identity deception scenarios.
@@ 3.4. Dataset Statistics
We note that existing datasets (e.g., HDTF and CelebV-HQ) may exhibit gender and ethnic biases. To address this, we prioritize diversity across four dimensions during video selection(Fig. 2): (1) Gender balance: We achieve a 57:43 male-female ratio to reflect real-world demographics and avoid excessive gender skew. (2) Ethnic diversity: The dataset includes White, African American, and Asian individuals, ensuring broad racial representation and enhancing model generalization. (3) Nationality and cultural background: Subjects from diverse countries and cultural contexts are selected to reduce regional bias and enhance global representativeness. (4) Scenario diversity: Video content simulates real-world scenarios, including social interactions, news broadcasts, and daily conversations, to improve applicability in practical contexts.
@@ 3.5. User Study
To evaluate the quality of the DigiFakeAV dataset, we recruited 100 participants with expertise in computer vision research, as we deemed their experience and domain knowledge essential for effectively assessing forged videos. The user study covers DigiFakeAV and three prior datasets—FF-DF, Celeb-DF, and DFDC—representing the first, second, and third-generation deepfake video benchmarks, respectively.
We select 50 video clips from each dataset and provide participants with a platform to evaluate each video’s authenticity. Participants are asked to vote on whether each clip was real or fake. The results of the user study, summarized in Tab.  3, clearly demonstrate that the proposed dataset was rated as highly challenging, highlighting the realism and detection difficulty of synthetic videos in DigiFakeAV.
Table 3. User Ratings of Fake Video Authenticity Across Datasets.
Dataset FF-DF Celeb-DF DFDC DigiFakeAV (ours) Fake / Real (%) 12% 46% 32% 68%
4. DigiShield
We propose DigiShield, a multimodal audio-visual spatiotemporal fusion detector. Our approach focuses on identifying subtle variations in facial features over time in synthetic videos. To capture such fine-grained temporal inconsistencies, we employ 3D convolution to model spatial and temporal patterns jointly. The integration of audio further enhances robustness by providing synchronization cues that improve forgery localization.
@@ 4.1. Model Framework
We find that single-modality analysis is insufficient for detecting digital human forgery artifacts. Thus, it is crucial to jointly model spatial and temporal inconsistencies without over-relying on any single cue. Building on prior 3D CNN and two-stream approaches (wang2023altfreezinggeneralvideoface, ; masi2020twobranchrecurrentnetworkisolating, ), we further exploit audio-visual cues through multimodal fusion. Our method includes (1) a spatiotemporal two-stream pipeline, (2) audio-visual fusion, and (3) a classification layer (see Fig.4). The model is trained using both contrastive and cross-entropy losses to capture audio-visual inconsistencies indicative of forgery.
Spatiotemporal two-stream pipeline. Formally, given a set of video clips $x_{\raisebox{-0.90417pt}{\scriptsize in}}^{\raisebox{0.60275pt}{\scriptsize v}}\in\mathbb{R}^{T\times C\times H_{v}\times W_{v}}$ and audio signals $x_{\raisebox{-0.90417pt}{\scriptsize in}}^{\raisebox{0.60275pt}{\scriptsize a}}\in\mathbb{R}^{C\times H_{a}\times W_{a}}$ (where T denotes the number of frames in the same video, C represents the number of channels, and H × W indicates the spatial dimensions), we first employ a spatiotemporal two-stream pipeline based on 2D and 3D residual convolutions to capture forgery traces. The spatiotemporal representations of the video and audio can then be computed as follows:
@@
The variables $T^{\prime}$,$C^{\prime}$,$H^{\prime}$, and $W^{\prime}$denote the temporal, channel, height, and width dimensions, respectively, after processing through the convolution.
Multimodal spatiotemporal fusion. The spatiotemporal features of video and audio, denoted as $\mathbf{f}_{v}\in\mathbb{R}^{N_{v}\times d}$ and $\mathbf{f}_{a}\in\mathbb{R}^{N_{a}\times d}$ respectively, are extracted through the video encoder $\mathcal{F}_{v}$ and audio encoder $\mathcal{F}_{a}$, and subsequently flattened along their spatiotemporal dimensions. Here, $N_{v}$ and $N_{a}$ denote the number of spatiotemporal tokens for video and audio, respectively, and $d$ represents the feature dimension.The flattened features are denoted as $\widetilde{\mathbf{f}}_{v}\in\mathbb{R}^{(N_{v}d)}$ and $\widetilde{\mathbf{f}}_{a}\in\mathbb{R}^{(N_{a}d)}$, which serve as input tensors to the multi-head attention mechanism.
We then apply cross-attention and self-attention modules to enhance multimodal interaction and feature fusion as follows:
@@
Final decision layer. The fused feature obtained is mapped to a binary classification probability space to make the final decision on the input video.
Figure 4. Overview of the DigiShield model: it detects forgery traces by integrating spatiotemporal information from both video and audio modalities.
@@ 4.2. Loss Function
To effectively fuse visual and audio modal features, we adopt a contrastive loss defined as:
(3) $\small\mathcal{L}_{con}=\frac{1}{N}\sum_{i=1}^{N}\left[y_{i}\,D(\widetilde{\mathbf{h}}_{v}^{(i)},\widetilde{\mathbf{h}}_{a}^{(i)})^{2}+(1-y_{i})\,\max\big{(}0,m-D(\widetilde{\mathbf{h}}_{v}^{(i)},\widetilde{\mathbf{h}}_{a}^{(i)})\big{)}^{2}\right]$
where $N$ is the batch size, $\widetilde{\mathbf{h}}_{v}^{(i)}$ and $\widetilde{\mathbf{h}}_{a}^{(i)}$ are the flattened visual and audio features of the $i$-th sample, $y_{i}\in\{0,1\}$ denotes ground-truth labels (1 for genuine, 0 for fake), $D(\cdot,\cdot)$ is the Euclidean distance, i.e., $D(\mathbf{x},\mathbf{y})=\|\mathbf{x}-\mathbf{y}\|_{2}$, and $m>0$ is a margin parameter.
@@
Here, ${L}_{con}$ and $\mathcal{L}_{ce}$ denote the contrastive loss and the cross-entropy loss, respectively.
@@ 5.1. Implementation Details
Preprocessing. During the preprocessing stage, we perform specialized processing on video and audio data. For the video modality, we use the RetinaFace face detection model to locate facial regions and perform corresponding cropping. For the audio modality, we first extract audio from videos, resample it to 16 kHz, and store it in WAV format. We then compute the Mel-frequency cepstral coefficients (MFCC) for each audio frame and convert them into a three-channel image representation as the audio input for the multimodal model.
Evaluation Protocol. For DigiShield, we adopt 3D-ResNet-50 as the default backbone. To ensure a fair and comprehensive evaluation of our proposed dataset, all experiments are conducted on data splits that follow an 8:1:1 ratio for training, validation, and testing, respectively, with no identity overlap between training and testing subsets. During training, we uniformly sample 20 frames from DF-TIMIT and 30 frames from DigiFakeAV videos, and apply data augmentation including color jittering and random cropping. To address the class imbalance in DigiFakeAV, we further apply heuristic upsampling to real samples during training, ensuring a balanced distribution between real and fake data. The AUC was employed as the primary evaluation metric. All experiments were conducted under consistent data preprocessing and training conditions to guarantee the stability and reliability of the findings.
Table 4. AUC scores (%) of various methods on compared datasets.
@@
5.2. Baselines Comparison
To conduct a thorough evaluation of the DigiFakeAV dataset, we systematically select nine representative deepfake detection algorithms based on the criteria of representativeness and accessibility. These methods encompass six distinct categories: Basic (Naive), Spatial, Temporal, Frequency, Hybrid Domain, and Multi-modal approaches. The chosen models comprise Meso4 (Afchar_2018, ), MesoInception4 (Afchar_2018, ), Xception-c23 (rössler2019faceforensicslearningdetectmanipulated, ), Capsule (nguyen2019usecapsulenetworkdetect, ), HeadPose (yang2018exposingdeepfakesusing, ), F3-Net (qian2020thinkingfrequencyfaceforgery, ), Cross Efficient ViT (Coccomini_2022, ), SFIConv (10286083, ), and SSVF (feng2023selfsupervisedvideoforensicsaudiovisual, ).
We follow standard benchmarking protocols, with the precision setting for baseline experiments rounding all classification scores to three decimal places ( $\ 10^{-3}$). As shown in Tab. 4, most evaluated methods exhibit significant performance degradation on DigiFakeAV compared to their results on other public datasets. For example, the AUC of SFIConv drops from 100.0% (DF-TIMIT HQ) to 71.2%, and SSVF declines from 94.5% (FakeAVCeleb) to 51.0%. These results highlight the effectiveness of digital human generation technologies in evading detection and expose the vulnerability of current models to advanced forgery strategies. A comprehensive description of the benchmark methods employed and further experimental evaluations are provided in the Supplementary Materials for reference.
@@ 5.3. DigiShield Comparison
As shown in Tab.  5, our model achieves an AUC of 80.1 on the more challenging DigiFakeAV dataset, outperforming existing methods and indicating that the digital human forgeries we propose require richer multimodal cues for effective detection. Our approach also achieves state-of-the-art (SOTA) performance on the DF-TIMIT dataset with an AUC score of 1.0, demonstrating its strong generalization ability across data of varying qualities.
Table 5. Comparison of AUC scores for DigiShield on DF-TIMIT and DigiFakeAV.
Method DF-TIMIT-LQ DF-TIMIT-HQ DigiFakeAV
MesoInception4 80.4 62.7 63.8
Capsule 78.4 74.4 65.3
Xception-c23 95.9 94.4 66.1
F3-Net 99.8 99.4 66.4
SFIConv 100.0 100.0 71.2
DigiShield 100.0 100.0 80.1
Table 6. Ablation study of AVTSF in DigiFakeAV.
@@ 5.4. Ablation Study
Tab. 6 validates the effectiveness of our audio-visual spatiotemporal fusion strategy. Using only cross-entropy loss ($L_{ce}$) achieves an AUC of 73.6, while adding contrastive loss ($L_{con}$) and Cross-attention (CrossAtt) improves it to 77.4. Self-attention (SelfAtt) further enhances temporal coherence, achieving an AUC of 80.1.
These results highlight the importance of both inter-modal and intra-modal attention mechanisms in detecting subtle digital human artifacts and diffusion-based forgery. For more detailed information, please refer to the Supplementary Materials.

### v2

@@ 3.3. Dataset Organization
Forged videos and audio can form four combinations: (1) Real Video–Real Audio (RV-RA), (2) Real Video–Fake Audio (RV-FA), (3) Fake Video–Real Audio (FV-RA), and (4) Fake Video–Fake Audio (FV-FA). We exclude RV-FA for two reasons: first, we focus on video-based facial forgery detection, whereas RV-FA does not involve human facial manipulation. For example, DFDC (dolhansky2020deepfakedetectionchallengedfdc, ) labels RV-FA samples as fake, which is inherently unfair to vision-only detectors. Second, audio-only manipulations without visual changes are often obvious and pose limited challenges for detection. For instance, FakeAVCeleb (khalid2022fakeavcelebnovelaudiovideomultimodal, ) has audio-video synchronization delays, which are relatively easy to detect.
We then introduce three combinations and the corresponding methods adopted, with detailed explanations:
@@
Fake Video–Real Audio (FV-RA) This category contains fake videos generated by real audio-driven diffusion models. We first extract and convert audio from real videos into WAV format. RetinaFace (deng2019retinafacesinglestagedenseface, ) is then used to detect and crop representative facial frames. Using five digital human generation methods: Sonic (ji2024sonicshiftingfocusglobal, ), Hallo1 (xu2024hallohierarchicalaudiodrivenvisual, ), Hallo2 (cui2024hallo2longdurationhighresolutionaudiodriven, ), Echomimic (chen2024echomimiclifelikeaudiodrivenportrait, ), and V-Express (wang2024vexpressconditionaldropoutprogressive, ), we generate 25,000 forged videos.
Fake Video–Fake Audio (FV-FA). This category encompasses both forged audio and video. We employ a SOTA voice cloning method, CosyVoice 2 (du2024cosyvoice2scalablestreaming, ). Specifically, we first use a large language model to generate processed text. By providing real audio-text pairs, the CosyVoice 2 model synthesizes fake audio mimicking the target speaker’s voice characteristics. We then generate 25,000 forged videos based on this fake audio using algorithms like Sonic, Hallo1, Hallo2, and Echomimic.
@@ 3.4. Dataset Statistics
Previous datasets such as HDTF (zhang2021flow, ) and CelebV-HQ (zhu2022celebvhqlargescalevideofacial, ) exhibit demographic imbalances in data distributions. For example, HDTF contains a male-to-female ratio of 65:35, while CelebV-HQ includes predominantly Caucasian subjects. To address this, we show in Fig. 2 how we prioritize diversity across four dimensions during video selection: (1) Gender balance: we achieve a 57:43 male-to-female ratio to reflect real-world demographics and avoid excessive gender skew. (2) Ethnic diversity: the dataset includes White, African American, and Asian individuals, ensuring broad racial representation and enhancing model generalization. (3) Nationality and cultural background: subjects from diverse countries and cultural contexts are selected to reduce regional bias and enhance global representativeness. (4) Scenario diversity: video content simulates real-world scenarios, including social interactions, news broadcasts, and daily conversations, to improve applicability in practical scenarios.
@@ 3.5. User Study
To evaluate DigiFakeAV, we recruit 30 computer vision researchers, leveraging their domain expertise to assess forged videos. The study includes DigiFakeAV and three established datasets—FF++ (rössler2019faceforensicslearningdetectmanipulated, ), Celeb-DF (li2020celebdflargescalechallengingdataset, ), and DFDC (dolhansky2020deepfakedetectionchallengedfdc, ), which represent the first to third generations of deepfake video benchmarks.
We randomly select 50 video clips from each dataset and present them to participants via a dedicated evaluation interface. Participants are asked to judge each clip’s authenticity by voting whether it is real or fake.
As summarized in Tab. 3, the proposed DigiFakeAV dataset yields the highest misclassification rate among all tested datasets, with 68% of videos erroneously classified, comprising both ”real” (fake videos are classified as real) and ”fake” (real videos are classified as fake).
Table 3. Misclassification rates of videos labeled as ”fake” or ”real” in the user study across deepfake datasets.
Dataset FF++ Celeb-DF DFDC DigiFakeAV (ours) ”Fake” or ”real” (%) 12% 46% 32% 68%
4. DigiShield: A Multimodel Deepfake Detection Baseline
Existing detection models show substantial performance degradation on the DigiFakeAV dataset (detailed in the Sec. 5.2, Baseline Comparison), highlighting their limitations in capturing multimodal features of diffusion-based digital human forgeries. To address this, we introduce DigiShield, which enhances adaptability to novel forgery types and serves as a reproducible baseline and reference for future research.
@@ 4.1. Model Framework
Single-modality analysis proves inadequate for digital human forgery detection, making it essential to jointly model spatial and temporal inconsistencies. Building on prior spatiotemporal and two-stream networks (wang2023altfreezinggeneralvideoface, ; masi2020twobranchrecurrentnetworkisolating, ), we incorporate multimodal audio-visual cues. As shown in Fig. 4, our method includes (1) a spatiotemporal two-stream pipeline, (2) audio-visual fusion, and (3) a classification layer.
Spatiotemporal two-stream pipeline. Formally, given a set of video clips $x_{\raisebox{-0.90417pt}{\scriptsize in}}^{\raisebox{0.60275pt}{\scriptsize v}}\in\mathbb{R}^{T\times C\times H_{v}\times W_{v}}$ and audio signals $x_{\raisebox{-0.90417pt}{\scriptsize in}}^{\raisebox{0.60275pt}{\scriptsize a}}\in\mathbb{R}^{C\times H_{a}\times W_{a}}$ (where $T$ denotes the number of frames in the same video, $C$ represents the number of channels, $H_{v}$ $W_{v}$ indicates the spatial dimensions, The Mel-spectrogram $H_{a}\times W_{a}$ represents the frequency bins ($H_{a}$) and time frames ($W_{a}$) dimensions), we first employ a spatiotemporal two-stream pipeline based on 3D and 2D residual convolutions to capture forgery traces. The spatiotemporal representations of the video and audio can then be computed as follows:
@@
where $\mathcal{F}_{v}$ and $\mathcal{F}_{a}$ denote the visual and audio encoders respectively, the variables $T^{\prime}$, $C^{\prime}$, $H^{\prime}_{v}\times W^{\prime}_{v}$, and $H^{\prime}_{a}\times W^{\prime}_{a}$ denote the temporal, channel, spatial dimensions, and frequency-time representation of the video and audio features after convolutional and pooling operations, respectively.
Multimodal spatiotemporal fusion. The classification layer of visual and audio encoders is redesigned to produce spatiotemporal features for cross-modal attention fusion. Specifically, the extracted feature maps from both the video and audio streams, denoted as $f_{v}$ and $f_{a}$, are reshaped into flattened representations: $\widetilde{f}_{v}\in\mathbb{R}^{N_{v}\times d}$ and $\widetilde{f}_{a}\in\mathbb{R}^{N_{a}\times d},$ where $N_{v}=T^{\prime}\times H^{\prime}_{v}\times W^{\prime}_{v}$ and $N_{a}=H^{\prime}_{a}\times W^{\prime}_{a}$, with $d=C^{\prime}$ indicating the dimension of the feature channel.
We then employ cross-attention and self-attention modules to enhance multimodal interaction and feature fusion as follows:
@@
Final decision layer. After extracting audio and video features, the fused representation is formed by concatenating these multimodal features along the channel dimension, effectively integrating complementary information across modalities. These combined features are then processed through a series of downsampling (Down) modules to progressively reduce spatial redundancy while preserving salient characteristics. Finally, the refined representation is fed into a fully connected (FC) layer, which maps the feature embedding into a binary classification probability space, facilitating the final decision on the authenticity of the input video.
Figure 4. Overview of the DigiShield, it detects forgery traces by integrating spatiotemporal information from both video and audio modalities.
[…]
Overview of the DigiShield
@@ 4.2. Loss Function
The model is trained using both contrastive and cross-entropy losses to capture audio-visual inconsistencies indicative of forgery. To effectively fuse visual and audio modal features, we adopt a contrastive loss defined as:
(3) $\mathcal{L}_{con}=\frac{1}{N}\sum_{i=1}^{N}\big{[}$ $\ y_{i}\,D(\widetilde{\mathbf{h}}_{v}^{(i)},\widetilde{\mathbf{h}}_{a}^{(i)})^{2}$
$+(1-y_{i})\,\max\big{(}0,m-D(\widetilde{\mathbf{h}}_{v}^{(i)},\widetilde{\mathbf{h}}_{a}^{(i)})\big{)}^{2}\big{]}$
where $N$ is the batch size, $\widetilde{\mathbf{h}}_{v}^{(i)}$ and $\widetilde{\mathbf{h}}_{a}^{(i)}$ are the flattened visual and audio features of the $i$-th sample, $y_{i}\in\{0,1\}$ denotes ground-truth labels (1 for real, 0 for fake), $D(\cdot,\cdot)$ is the Euclidean distance, e.g., $D(\mathbf{x},\mathbf{y})=\|\mathbf{x}-\mathbf{y}\|_{2}$, and $m>0$ is a margin parameter.
@@
This composite loss serves as the training objective.
@@ 5.1. Implementation Details
During preprocessing, videos are processed with RetinaFace (deng2019retinafacesinglestagedenseface, ) for facial detection and cropping, while audio tracks are extracted, resampled to 16 kHz, and saved as WAV files. Mel-frequency cepstral coefficients (MFCCs) are computed for each frame and transformed into three-channel images as audio inputs for the multimodal model.
For DigiShield, we use 3D-ResNet-50 as the backbone. On DigiFakeAV [^4] Details of training setting are in the Appendix C, experiments adopt an 8:1:1 train/val/test split with no identity overlap, sampling 30 frames per video and applying color jitter and random cropping for augmentation. Heuristic upsampling is used to balance real and fake samples. The Area Under the Curve (AUC) is the primary evaluation metric, with all preprocessing and training settings kept consistent for fair comparison. We also use DF-TIMIT (korshunov2018deepfakesnewthreatface, ) and process the dataset following the official settings.
Table 4. AUC scores (%) of various methods on previous datasets and our DigiFakeAV.
@@
5.2. Baseline Comparison
To conduct a thorough evaluation of the DigiFakeAV dataset, we systematically select nine representative deepfake detection algorithms based on the criteria of representativeness and accessibility. The selected models cover six categories: Naive (Meso4 (Afchar_2018, ), MesoInception4 (Afchar_2018, ), Xception-c23 (rössler2019faceforensicslearningdetectmanipulated, )), Spatial (Capsule (nguyen2019usecapsulenetworkdetect, ), HeadPose (yang2018exposingdeepfakesusing, )), Frequency (F3-Net (qian2020thinkingfrequencyfaceforgery, )), Temporal (Cross Efficient ViT (Coccomini_2022, )), Hybrid Domain (SFIConv (10286083, )), and Multi-modal (SSVF (feng2023selfsupervisedvideoforensicsaudiovisual, )).
As shown in Tab. 4, existing deepfake detection methods universally face performance collapse on the DigiFakeAV dataset. For example, the AUC of SFIConv drops from 100.0% (DF-TIMIT HQ) to 71.2%, and SSVF declines from 94.5% (FakeAVCeleb) to 51.0%. These results validate the disruptive challenge posed by diffusion-based digital human forgeries to existing detection paradigms. Traditional methods relying on local texture anomalies or static frequency-domain features are insufficient to handle highly coherent multimodal synthetic content, urgently requiring a reconstruction of the detection of spatiotemporal dynamics and cross-modal interactions [^5] Further evaluations of baseline comparison are provided in the Appendix D.
@@ 5.3. DigiShield Comparison
As shown in Tab.  5, DigiShield achieves an AUC of 80.1% on the highly challenging DigiFakeAV dataset, representing a 57.1% and 12.5% improvement over the current best multimodal detector SSVF (51.0%) and frequency-domain method SFIConv (71.2%), respectively. This performance breakthrough stems from the model’s targeted modeling of the intrinsic features of diffusion-based digital human forgeries. Meanwhile, DigiShield maintains a 100% AUC on the traditional face-swapping dataset DF-TIMIT, validating its generalization. Further analysis indicates that DigiShield’s success confirms the effectiveness of multimodal spatiotemporal inconsistency as a core metric for next-generation detection.
Table 5. Comparison of AUC scores for DigiShield on DigiFakeAV and DF-TIMIT
Method DigiFakeAV DF-TIMIT-LQ DF-TIMIT-HQ
MesoInception4 63.8 80.4 62.7
Capsule 65.3 78.4 74.4
Xception-c23 66.1 95.9 94.4
F3-Net 66.4 99.8 99.4
SFIConv 71.2 100.0 100.0
DigiShield (ours) 80.1 100.0 100.0
Table 6. Ablation study of DigiShield in DigiFakeAV.
@@ 5.4. Ablation Study
Tab. 6 validates the effectiveness of our audio-visual spatiotemporal fusion strategy, we progressively introduce each component to assess its contribution. The baseline model uses a single-stream setting—by removing the audio extractor and training only with visual features under cross-entropy loss ($L_{ce}$)—which yields an AUC of 73.6%. This indicates that relying solely on single-modality classification is insufficient to capture inconsistency features. When the audio stream and contrastive loss ($L_{con}$) are incorporated to form a dual-stream architecture, and cross-modal attention (CrossAtt) is further introduced to explicitly model inter-modal discrepancies such as speech-lip asynchrony, the AUC is boosted to 77.4%. Finally, adding self-attention (SelfAtt) enhances long-term temporal coherence, pushing AUC to 80.1%. These results underscore the importance of both temporal coherence and cross-modal alignment in detecting digital human forgeries [^6] More details of the ablation study are in the Supplementary Materials (Appendix E).